OpenStudy (anonymous):

find the first three output values of the fractal generating function: f(z)=z^2+2+i i= imaginary number square root of -1

OpenStudy (anonymous):

Make sure you understand that (a+b)^2=a^2+2ab+b^2, and that i^2=-1. Then, just substitute and solve, plugging in the output value back into the function each time. (2+i)^2+2+i =4+4i-1+2+i=5+5i (5+5i)^2+2+i =25(1+i)^2+2+i=25(1+2i-1)+2+i=51i+2

OpenStudy (anonymous):

2+i, 5+5i, 2+51i answer